# Shape-specific fluctuations of an active colloidal interface

### v1

@@ IV.3 Orientational fluctuations
Figure 5: Scaling of height fluctuations in locally flat region. A $W_{h}$ for different locally flat lengths $N^{\prime}$, with the dynamic $\beta_{h}^{\prime}\approx 0.25$ indicated via dashed black line. B Roughness scaling for $W_{h}^{\text{SS}}(N^{\prime})$. Here, $N=312$ and $\tau_{f}=0.2$, $\Lambda=0.2$, $\tilde{\Lambda}=0.01$.
IV.4 Locally flat regimes
@@
In the study of fluctuations of circular interfaces, the fluctuations are typically measured about a locally flat region along the growing front [31, 32, 9]; and the roughness scalings thus correspond to those of a locally flat region. We repeat the same procedure here, and vary $N^{\prime}$ under the constraint of $N^{\prime}<<N$. In this section, $\frac{N^{\prime}}{N}<0.18$ such that the interface is effectively flat. The results are presented in Fig. 5. Probing the locally flat regime, we find that the height fluctuations in the early-time II and steady-state regimes, respectively, can be collapsed onto a FV-type scaling with growth exponent $\beta_{h}^{\prime}\approx 0.25$, and roughness exponent $\alpha_{h}^{\prime}\approx 1$. This would imply a dynamic exponent $z_{h}^{\prime}=4$. However, it is to be noted that this implies essentially neglecting the super-ballistic regime, where the data does not collapse. A dynamic exponent $z=4$ is reminiscent of driven interfaces with conservation laws [17, 36, 37], although the individual growth and roughness exponents differ. An appropriate continuum description of our system remains an open problem. The $W_{\theta}$ scaling in this regime is presented in Appendix .5.
Quantity $\beta_{h}^{(1)}$ $\beta_{h}^{(2)}$ $\alpha_{h}$ $z_{h}^{(1)}=\frac{\alpha_{h}}{\beta_{h}^{(1)}}$ $z_{h}^{(2)}=\frac{\alpha_{h}}{\beta_{h}^{(2)}}$
$W_{h}(N)$ $0.25(01)$ $1.53(02)$ $0.90(03)$ $3.61(25)$ $0.58(03)$
$W_{h}(N^{\prime})$ $0.25(01)$ NA $1.01(01)$ $4.04(20)$ NA
Table 1: Summary of exponents for $W_{h}$. Note that $\beta_{h}^{(1)}$ and $\beta_{h}^{(2)}$ refer to two sequential regimes for each before reaching saturation.
Quantity $\beta_{\theta}$ $\alpha_{\theta}$
$W_{\theta}(N)$ $1.00(00)$ 0.47(03)-0.47(03)
Table 2: Summary of exponents for $W_{\theta}$.
V Discussion and summary
@@
Roto-translational coupling has been well-studied for their various implications on dynamical and collective behaviour in active matter physics. The effect of this coupling on scaling laws have been established for the mean-squared displacement - i.e enhanced or anomalous diffusion - in various models (see for instance [38, 39, 40]). Here, we instead present the effect of roto-translational induced global topological change, and the resultant novel scaling laws for interface fluctuations. Remarkably, we find that the height fluctuations show an FV data collapse, with a new set of dynamic and roughness exponents; with the former rationalized via a simple mean-field analysis of roto-translational coupling. The orientational fluctuations, in addition, display novel scaling behaviour, notably smoothening with system size, which we attribute to the enhanced orientational rigidity of longer chains. The exponents obtained in this paper for height fluctuations are summarized in Table 1, while the exponents for orientation fluctuations are summarized in Table 2.
In equilibrium (passive) systems, the typical time scales that govern the height fluctuations are predominantly drift and diffusion (in addition to any deterministic interactions and hydrodynamics). In certain models one might in addition have an additional (Arrhenius) time scale that sets a time scale of local equilibration in an energy landscape [16, 41, 19]. The self-propelled, roto-translational colloidal chain instead possesses an additional timescale within which deterministic spatio-temporal (i.e topological) changes in the interface take place, leading to super-ballistic growth and smoothening phenomena.
We note that existing models of active interfaces have also reported additional time scales that arise due to non-equilibrium effects; for example see [19, 42, 43] that discuss proteins hopping on a membrane. Indeed, these works also report FV scaling laws in their systems, but this is (naively) to be expected as the continuum evolution of the height fields resembles that of a modified KPZ-like model. The FV scaling laws that we report, however, arise from a shape change of the interface, with the C-shape topology being distinct from either a circular or flat interface.
Potential open questions thus naturally arise as to what appropriate (tractable) coarse-grained (continuum) models could describe the active interface presented here. It remains to be seen whether or not both the C-shape and locally flat topologies could be explained by the same continuum model. In addition, instead of periodic boundary conditions typically used to model growing interfaces, “clamped” boundary conditions are required for the curvature to set in. These ideas provide challenges for further work.
To test the predictions here, we note that the experimental paradigm requires long-ranged internal repulsions within the monomers in the chain, that give it its rigid interface structure [24]. To observe significant fluctuations in these systems, if one equates damping forces to thermal forces, $\eta bv_{s}\sim k_{b}T/b$ ($\eta$ the viscosity of the medium and $v_{s}$ the self-propulsion $\approx 10\mu ms^{-1}$), we obtain $b\approx 10^{-7}m$ (at least 10 times smaller than typical chemically interacting colloidal systems). Naively, these would correspond to active polymer regimes of biological interest [44, 45]. However, recent findings [46] have reported how phase change in 8CB liquid crystal emulsions can be induced via external temperature changes. In the nematic/isotropic phases, the individual colloids no longer deterministically self-propel but in addition have a fluctuating propulsion direction. This would correspond to finite $\tau_{r}$ in our model. In addition, these results could also be relevant to active interfacial systems such as colonies of migrating bacteria [47], where chemical interactions are typically long-ranged. Another open problem, would be the extension of this model to a ring topology, where the possibility of attaining spin waves is imminent [21].
@@
.1 Early time approximation and probability distribution
@@
These height fluctuations are alternatively characterized by the distribution of fluctuations, $P(W^{2})$ [32, 48]. For the typical EW interface (non-propelling), known results exist and are re-written here. We can directly quote the result from [48] for arbitrary initial height distribution
$\langle W^{2}\rangle_{SS}P(W^{2},t)=\int_{-i\infty}^{i\infty}\frac{dy}{2\pi i}e^{xy}\prod_{n=1}^{\infty}\frac{-ys^{2}_{n0}e^{-n^{2}\tau}/(1+ya_{n})}{1+ya_{n}},$ (21)
@@
$x=\frac{W^{2}}{W^{2}_{SS}},~\tau_{p}=\frac{8\pi^{2}\nu_{h}t}{L^{2}},~s_{n0}^{2}=\frac{2|c_{n}(0)|^{2}}{W^{2}_{SS}}\,.$ (22)
For a flat initial height profile, one has $s_{n0}=0$, which gives [48]
$\langle W^{2}\rangle_{SS}P(W^{2},t)=\sum_{m=1}^{\infty}\frac{1}{a_{m}}\exp\left[\frac{-x}{a_{m}}\right]\prod_{n=1,\neq m}^{\infty}\frac{a_{m}}{a_{m}-a_{n}}\,.$ (23)
The terms $c_{n}(0)$ are Fourier transforms of the initial height profile. In our case, we start with $h(x,t=0)=0$, which fixes $c_{n}(0)=0$.
@@
The early time approx of (20) corresponds to setting $\Lambda\rightarrow\infty$. Note that one can either $\tau<<\tau_{f}$ probe or $\chi_{r}=0$ globally; we pick the latter. The analytical distribution of (23) can then be compared to this limit; we find that the distribution is reproduced - black line and blue scatter of Fig. 6(A). We can thus suitably compare this with the case of $\Lambda$ finite, restricting to $\tau<<\tau_{f}$. Here, we find that our propelled interface displays deviations from this passive distribution; this is displayed in Fig. 6(B) - yellow scatter. For both cases, we plot alongside the best fit analytical distribution in dashed black, and we further find that the averaged least squared error of the fit is substantially larger in the case of $\Lambda$ finite; we thus conclude that the early time non-equilibrium distribution deviates from the passive EW case. This marginal deviation at early times arises again via the aforementioned roto-translational coupling; this thus again constitute a non-equilibrium signature that appears via deterministic (but small) orientational interactions. This deviation in the height distribution is analogous to that seen in so-called “anomalous diffusion” [49, 50], where microscopic colloidal particles in specific systems display a mean-squared displacement exponent of $1$, but the probability distribution of displacements has been found to follow the Laplace (instead of Gaussian) distribution [50, 51]. The exact analytical distribution of the chemically interacting chain is so-far intractable, and thus remains an open future problem.
.2 Mean-field approximation of super-ballistic growth exponent
In the main text we have reported the growth exponent of the CI to be $\beta_{h}\approx 1.5$ (super-ballistic). Although rationalizing this for a curved surface is beyond exact analytical calculation (in this work); we may study the flat interface (thermodynamic limit), via a mean-field analysis.
Let us study the simplest case of roto-translational coupling, in particular where (small) angular dynamics set in. Consider the $i$th positional dynamics from (2). For small $\theta_{i}$, we can write this as
@@
where $S_{i}=\frac{1}{4b^{2}}\sum_{j}{\frac{1}{(i-j)^{2}}\Big{[}2\Theta(i-j)-1\Big{]}}$, with $\Theta(x)$ the step function. The deterministic solution to this is
@@
We next use the steady-state approximation $\frac{t}{\tau_{f}}<<\frac{1}{b^{2}S_{i}}$. Thus, in the growth regime of $t\sim\tau_{f}$, this requires $b^{2}S_{i}>>1$, which is satisfied in the thermodynamic limit.
We then plug (26) into (24). We then use the small argument expansion for $\cosh(z)\approx 1+\frac{z^{2}}{2}$, with $z=\frac{\chi_{r}S_{i}t}{2}$. This gives us
$\dot{x}_{i}=F_{i}(t)\approx v_{s}\Big{(}1-\frac{b^{4}}{2\tau_{f}^{2}}S_{i}^{2}t^{2}\Big{)}+F^{b}_{x,i}+\xi_{t,i}$ (27)
@@
$\langle X\rangle=v_{s}t-\frac{b^{4}}{6\tau_{f}^{2}}\langle S^{2}\rangle t^{3}$ (28)
where $\langle\rangle$ is performed over both the monomers and realizations. Here, $X$ is a site-independent average of the chain position. We obtain the scaling for $(\Delta h)^{2}$ as follows
$W_{h}^{2}$ $\approx\frac{b^{4}}{6\tau_{f}^{2}}\langle\Big{(}S_{i}^{2}-\langle S\rangle^{2}\Big{)}^{2}\rangle t^{6}+2D_{t}t$ (29)
where we have used that $\langle F^{b}(t^{\prime})\rangle=0$ and $\langle F^{b}(t^{\prime})F^{b}(t^{\prime})\rangle=\langle F^{b}(t)\rangle\langle F^{b}(t^{\prime})\rangle$. Strictly speaking, the latter will cease to be true when curvature sets in (spring forces then become correlated), but it will suffice for the flat approximation used here. We thus obtain the result that the dynamics is super-ballistic with exponent $\sim 3$, when orientational contributions are minimally accounted for such that $\text{Var}(S)\neq 0$. We thus conclude that minimal roto-translational coupling on a flat interface can predict super-ballistic growth exponents but clearly overestimates that of the surface with finite curvature.
.3 Curvature of chain
@@
where $h_{i}$ is the height function of the chain, evaluated at each monomer location, measured from the vertical line connecting the edge monomers. The average is performed across the monomers in the chain, giving a scalar value. In our case, the inner bracket is simply reduced to $\partial_{y}\left[{\partial_{y}x_{i}(y)}/{\sqrt{1+|\partial_{y}x_{i}|^{2}}}\right]$. For Fig. 1, we take the averaged $|\kappa|$ in the steady state. Further, we note that $\kappa$ $\rightarrow 0$ as $N\rightarrow\infty$ [28]. For a circular interface, $\kappa$ would be scale as $1/R$ ($R$ the radius of the circle), whereas for a flat interface $\kappa$ would be constant. Thus, this variation $\kappa(N)$ profile can be used as proxy definition for the C-shape.
.4 Other DSS for large NN
It is to be noted also that the phase diagram in Fig. 1 being $N$ specific and only comparing $\Lambda$, $\tilde{\Lambda}$ does exclude other DSS that arise from simulation. Examples of these are shown in Fig. 7. Four examples are shown, we label them as (i) “Alt-C”, (ii) “Random front”, (iii) “Stiff-C”, and (iv) “Pinched front” respectively. We note that DSS(ii) corresponds to the “Disordered” phase in Fig. 1. For each of these DSS, the chain propels in a deterministic direction along the x-axis; thus (8) can be computed in principle. However, none of these DSS satisfy all the criteria enumerated in Sec. IV.1. For DSS(i)-(iv), criteria (i) 2, 3, and 4, (ii) 1 and 3, (iii) 1,3 and 4; and (iv) 2 and 3 are satisfied. Thus, the CI is the unique topology that satisfies all four criteria. The hypothetical fluctuations for $W_{\theta}$ for these can also in principle be computed, but we do not pursue it here. In general, whenever there is an orientational symmetry along the body axis (e.g. for DSS(i)), we expect a smoothness exponent.
@@
.5 Scaling of $W_{\theta}$ in the locally flat regime

### v2

@@ V.3 With finite rotational noise
Figure 5: Scaling of height fluctuations in locally flat region. A $W_{h}$ for different locally flat lengths $N^{\prime}$, with the dynamic $\beta_{h}^{\prime}\approx 0.25$ indicated via dashed black line, fit between $(t/\tau N^{\prime z_{h}^{\prime}})\in[5\times 10^{-6},10^{-5}]$. B Roughness scaling for $W_{h}^{\text{SS}}(N^{\prime})$. Here, $N=312$ and $\tau_{f}=0.2$, $\Lambda=0.2$, $\tilde{\Lambda}=0.01$.
V.4 Locally flat regimes
@@
In the study of fluctuations of circular interfaces, the fluctuations are typically measured about a locally flat region along the growing front [36, 37, 9]; and the roughness scalings thus correspond to those of a locally flat region. We repeat the same procedure here, and vary $N^{\prime}$ under the constraint of $N^{\prime}<<N$. In this section, $\frac{N^{\prime}}{N}<0.18$ such that the interface is effectively flat. The results are presented in Fig. 5. Probing the locally flat regime, we find that the height fluctuations in the early-time ($\beta^{(1)}$) and steady-state regimes, respectively, can be collapsed onto a FV-type scaling with growth exponent $\beta_{h}^{\prime}\approx 0.25$, and roughness exponent $\alpha_{h}^{\prime}\approx 1$. This would imply a dynamic exponent $z_{h}^{\prime}=4$. This combination of exponents is reminiscent of interfaces with conservation laws in two dimensions  [38, 17, 41, 42]. However, it is to be noted that this implies essentially neglecting the super-ballistic regime, where the data does not collapse. The $W_{\theta}$ scaling in this regime is presented in Appendix .6.
VI Summary and Discussions
@@
Roto-translational coupling has been well-studied for their various implications on dynamical and collective behaviour in active matter physics. The effect of this coupling on scaling laws have been established for the mean-squared displacement - i.e enhanced or anomalous diffusion - in various models (see for instance [33, 43, 44]). Here, we instead present the effect of roto-translation-induced global topological change, and the resultant novel scaling laws for interface fluctuations. Remarkably, we find that the height fluctuations show an FV data collapse, with a new set of dynamic and roughness exponents; with the former rationalized via a simple linearised analysis of roto-translational coupling. The orientational fluctuations, in addition, display novel scaling behaviour, notably smoothening with system size, which we attribute to the enhanced orientational rigidity of longer chains.
Some of our observed exponents are rationalizable via comparison with those of well-studied models. The roughness exponent 0.9~0.9 accompanying the height fluctuations is higher than that of non-conserving mass models ($\sim 0.5$) and lower than conserving mass models ($\sim 1.5$). These two correspond to the limit of “monomeric units” exchanging with the bulk (e.g. evaporation) and being distributed across the surface respectively. The latter limit for instance corresponds to the dynamic exponent we report in the locally flat regime (with monomeric redistribution but discounting the C-shape). We thus see why the C-shape roughness lies in between these two values - the C-shape does distribute monomeric displacements across the body (thus less suppression of large scale fluctuations), but does so whilst maintaining a fixed bending rigidity. The growth exponent here being super-ballistic is solely a consequence of topological change due to roto-translational coupling; a simple picture as to how roto-translational coupling was give rise to super-ballistic growth is provided in Appendix .3, where an exact solution to a linearized approximate dynamics is presented (a flat chain with lowest-order roto-translational coupling). For the orientational sector, the smoothness can be appreciated by noting that as the system size increases, a proportionately larger fraction of the monomers are polarized along the x-axis, thus orientational fluctuations at larger length scales get more suppressed. The ballistic growth exponent on the other hand arises due to the deterministic effect of orientational changes that are disproportionately on the chain edges.
The emergence of novel scaling exponents [Table 1], phases and transitions between them [Fig. 1] demand a thorough hydrodynamic analysis. While we defer the proposition and analysis of such a theory to future work, here we briefly outline some ingredients of a requisite continuum picture. Although individually each colloid can self-propel in any direction, the emergent macroscopic chain spontaneously breaks the isotropy by moving in a chosen direction. When the chain is in its stiff phase (I in Fig. 1), all monomers point in the same direction. However, the C-shape phase (IV in Fig. 1) arises as a result of anti-alignment of monomer orientations. At the minimal level, it may then be hypothesized that the one-dimensional chain can be modeled in terms of two dynamically coupled fields: (i) the local interface height ($h(x,t)$) and (ii) the local polarization (${\bf P}(x,t)=P(\cos\theta(x,t),\sin\theta(x,t))$). In particular, we expect the height field to undergo noisy advection-diffusion dynamics where the advection direction is determined by the $\bf P$} field. On the other hand, $\bf P$} should follow a dynamics that not only allows spontaneous symmetry breaking ($\partial_{t}{\bf P}\propto a{\bf P}+b|{\bf P}|^{2}{\bf P}$) but also contains alignment and anti-alignment terms ($\partial_{t}{\bf P}\propto D_{1}\nabla^{2}{\bf P}+D_{2}\nabla^{4}{\bf P}$). We expect the transition from the stiff to C-shape to be accompanied by a change in sign of the coefficient of the aligning term $D_{1}$. How does $h(x,t)$ affect the dynamics of $\bf P$}? For both phases I and IV of Fig. 1, the local orientation is along the local normal to the interface. In other words, the local orientation is driven by the local curvature ($\nabla^{2}h$) of the chain. In effect, we expect such curvature-driven polarization to then introduce non-local forces in the dynamics of the height field. Further, instead of periodic boundary conditions typically used to model growing interfaces, “clamped” boundary conditions are potentially required for the curvature to set in. These ideas provide challenges for further work.
It is of further relevance to note where these results - the existence of growth and roughness exponents that are unique to this particular topology - fit in the broader study of active interfaces/membranes. Firstly, it is noted that non-equilibrium extensions to passive (growing) interfaces models have been theoretically proposed and new universality classes proposed. For instance, studying the capillary fluctuations on a non-equilibrium field theory, it has been shown that (in 2D) $\alpha=1/2$ and $z=3$, whilst additional non-linear terms decreases these values slightly [45, 46, 23]. These interfaces arise from underlying scalar fields derived from conservation laws whose dynamics violate symmetries found in equilibrium. A parallel set of active membrane models are those where a Monge-like membrane is coupled to (the concentration field) a membrane-spanning protein pump that exerts a mechanical force on the membrane upon exchanging material with it (chemical coupling), with solutions to these exhibiting travelling waves [47, 48, 49]. A recent set of works study a variant of this where the protein dynamics are inclusions on the height field which hop along the membrane with an additional time scale [19, 50, 51]. For these sets of “active membrane” models, the height correlations are characterized by $S(q)=\langle h(\mathbf{q})h(\mathbf{-q})\rangle\sim 1/q^{2}$ universally, the result for a Monge (passive) interface with Gaussian fluctuations [52]. This gives rise to the standard $\alpha=1/2$ roughening. For orientational correlations, though these have not been studied elsewhere to our knowledge, perhaps the closest set of works pertain to those studying tangential angular correlations on a liquid–liquid interface undergoing phase separation, where the tangential angular correlations show a smoothening at large scales [21, 53]. It is evident that the physics of these aforementioned systems are starkly different from that of the autophoretic C-shape chain; we nevertheless mention them here for the sake of completion.
It remains an open challenge to verify these exponents in an experimental system. To observe significant positional fluctuations in experimental realizations, if one equates damping forces to thermal forces, $\eta bv_{s}\sim k_{b}T/b$ ($\eta$ the viscosity of the medium and $v_{s}$ the self-propulsion $\approx 10\mu ms^{-1}$) (e.g. for droplet experiments [25, 54, 55]), one obtains $b\approx 10^{-7}m$, thus around 10 times smaller than typical chemically interacting colloidal systems. On the other hand, other phoretic colloidal systems (typically electrophoretic and bubble propulsion mechanisms) have reported propulsion speeds of $v_{s}\sim 1\mu ms^{-1}$ [56, 57], where the thermal fluctuations will be appreciable. Constructing extended propelling objects with a well-defined propelling interface out of these units remain an open challenge. The smaller colloids of $b\sim 10^{-7}m$ deserve a specific consideration. These correspond to so-called “nanomotor” systems, sub-micron (mostly metallic) colloids whose synthesis routes are well established [58, 59] and have been shown to self-propel via various mechanisms [60, 61, 62, 63]. Noting that $\tau_{r}\propto b^{3}$, we find that with $b\sim 10^{-7}m$, $\tau_{r}\sim 10^{-2}s$, which we note is smaller than the regime required for C-shape formation to take place (c.f. Section V.3). Thus, it may be feasible to only observe the orientational growth and smoothening with an appreciable $b\sim 10^{-6}m$, where $\tau_{r}\sim 10^{1}s$, which supports the C-shape. At this scale it is also important to note that motion is no longer restricted to be in two dimensions, and continuum theories of phoretic/fluid interactions may in addition break down [64, 65]. In this context it is of interest to note recent works that have reported how phase change in 8CB liquid crystal emulsions (micron-sized) can be induced via external temperature changes [54]. In the nematic/isotropic phases, the individual colloids no longer merely deterministically self-propel but in addition have a fluctuating propulsion direction. This would correspond to finite $\tau_{r}$ in our model without sacrificing monomeric size. Creating extended objects out of such systems, and/or realizing a robust C-shape with positional fluctuations thus remain an open problem. Finally, though extended colloidal chains propelled via external (as opposed to autophoretic) mechanisms have widely been synthesized [66, 67, 68, 69, 26, 27], neither have them have reported well defined shapes where interfacial height deviations and fluctuations may be readily computed - i.e. neither of these reported shapes that propel perpendicular to the body axis [70, 29]. It is noted that these systems suffer from the same translational noise suppression and noise scaling issues mentioned above.
@@
.1 Incorporation of chemical trails
Figure 6: Fluctuations of model with chemical trails. Example of A $W_{h}$ and $W_{\theta}$ evolution, with the dynamic $\beta_{h}$ labelled (fit between $t/(\tau N^{z_{h}})\in[100,200]$) and $t/\tau\in[1,10]$ respectively. B Orientation profile of the C-shape, taken at different time points. C Dynamical evolution of $\theta_{i}$ for selected monomers ii.. Here, $D_{c}=1$, $t_{0}=1$, with $N=64$ and $T=10^{5}$.
We present the results for $W_{h}$ and $W_{\theta}$ in Fig. 6 (A). We find that the scalings of $W_{h}$ are reproduced. With the micellar trails, when one starts such a system with arbitrary initial conditions (e.g. random), the trails deposited via (19) give rise to a forward-backward symmetry breaking in the orientational response to the chemicals. Thus, the system breaks the initial symmetry imposed on it and further universally picks up the C-shape. In our system, we choose sufficiently random initial conditions such that this effect is seen; we then compute $\Delta h$ once the chain has a well-defined interface.
For this section only one system size, $N=64$, was studied, averaged over $10$ realizations. Note that the dynamical evolution via (19) requires at each time step a sum over all previous time steps. Thus, even for this system size (on a desktop computer with processor frequency of 2.40GHz and 16 cores), one realization takes $3.77\times 10^{5}$ seconds. Thus, a rigorous system-size determination of roughening/smoothening is prohibitive. With only one curve for $W$, obtaining $\alpha$ and $\beta$ via (12) (14) requires self-consistency. For $W_{h}$, we fix $\alpha\approx 0.9$; in this setting the best fit obtained was $\beta_{h}=1.41\pm 0.01$. We have further checked that super-ballistic $\beta_{h}$ is always obtained if the coarsening is in between evaporative and mass conserving limits (this is solely a consequence of the C-shape topology and its fixed bending rigidity, irrespective of dynamical details; see arguments in Section VI).
Further, for $W_{\theta}$, we find that fixing $\beta_{\theta}=1$ gives $\alpha_{\theta}\approx 0.2$. Thus, not a smoothening exponent, but nevertheless much less rough than the standard growth models. If one attempts to fit a negative $\alpha_{\theta}$ instead, one instead finds super-ballistic orientational growth (e.g. $\alpha_{\theta}\approx-0.1$ gives $\beta_{\theta}\approx 4$). Super-ballisticity could also be anticipated given that the time evolution contains higher order temporal derivatives, specifically for monomers near the edges (e.g. $i=9,50,55$ in Fig.6(C) versus $i=311$ of Fig.3(D)), thus any fluctuations about the mean will deviate from a simple ballistic scaling. On the other hand, the steady-state angular profile is no longer fully polarized along the propulsion direction (compare Fig.6(B) versus Fig.3(C)), thus $\alpha_{\theta}$ is expected to be higher than in the case without trails.
A more complete picture obviously will arise with analyses at various system sizes. However, it is safe to conclude that (i) super-ballistic positional growth; (ii) roughnening of positional fluctuations; and (iii) a roughening of orientational fluctuations that is significantly less that standard models (e.g. EW/KPZ and/or Gaussian membranes) is present when chemical trails are included.
.2 Early time approximation and probability distribution
@@
These height fluctuations are alternatively characterized by the distribution of fluctuations, $P(W^{2})$ [37, 72]. For the standard EW interface (non-propelling), known results exist and are re-written here. We can directly quote the result from [72] for arbitrary initial height distribution
$P(W^{2},t)=\int_{-i\infty}^{i\infty}\frac{dy}{2\pi i}e^{xy}\prod_{n=1}^{\infty}\frac{-ys^{2}_{n0}e^{-n^{2}\tau}/(1+ya_{n})}{\langle W^{2}\rangle_{SS}\left(1+ya_{n}\right)},$ (22)
@@
$x=\frac{W^{2}}{W^{2}_{SS}},~\tau_{p}=\frac{8\pi^{2}\nu_{h}t}{L^{2}},~s_{n0}^{2}=\frac{2|a_{n}(0)|^{2}}{W^{2}_{SS}}.$ (23)
The terms $a_{n}(0)$ are Fourier transforms of the initial height profile. In our case, we start with $h(x,t=0)=0$, which fixes $a_{n}(0)=0$. For a flat initial height profile, one has $s_{n0}=0$, which gives [72]:
$P(W^{2},t)=\sum_{m=1}^{\infty}\frac{1}{a_{m}}\frac{\exp\left[\frac{-x}{a_{m}}\right]}{\langle W^{2}\rangle_{SS}}\prod_{n=1,\neq m}^{\infty}\frac{a_{m}}{a_{m}-a_{n}}.$ (24)
@@
The early time approx of (21) corresponds to setting $\Lambda\rightarrow\infty$. Note that one can either $\tau<<\tau_{f}$ probe or $\chi_{r}=0$ globally; we pick the latter. The analytical distribution of (24) can then be compared to this limit; we find that the distribution is reproduced - black line and blue scatter of Fig. 7(A). We can thus suitably compare this with the case of $\Lambda$ finite, restricting to $\tau<<\tau_{f}$. Here, we find that our propelled interface displays deviations from this passive distribution; this is displayed in Fig. 7(B) - yellow scatter. For both cases, we plot alongside the best fit analytical distribution in dashed black, and we further find that the averaged least squared error of the fit is substantially larger in the case of $\Lambda$ finite; we thus conclude that the early time non-equilibrium distribution deviates from the passive EW case. This marginal deviation at early times arises again via the aforementioned roto-translational coupling; this thus again constitute a non-equilibrium signature that appears via deterministic (but small) orientational interactions. This deviation in the height distribution is analogous to that seen in so-called “anomalous diffusion” [73, 74], where microscopic colloidal particles in specific systems display a mean-squared displacement exponent of $1$, but the probability distribution of displacements has been found to follow the Laplace (instead of Gaussian) distribution [74, 75]. The exact analytical distribution of the chemically interacting chain thus remains an open future problem.
.3 Linearised approximation of super-ballistic growth exponent
In the main text we have reported the growth exponent of the C-shape to be $\beta_{h}\sim 1.7$ (super-ballistic). Although rationalizing this for a curved surface is beyond exact analytical calculation in this work; we study instead a flat interface, via a the solution to the linearised approximate dynamics.
Let us study the simplest case of roto-translational coupling, in particular where (small) angular dynamics set in; in this setting the full set of equations are solvable. Consider the $i$th positional dynamics from (2). For small $\theta_{i}$, we can write this as
@@
where $S_{i}=\frac{1}{4b^{3}}\sum_{j}{\frac{1}{(i-j)^{2}}\Big[2\Theta(i-j)-1\Big]}$ for a completely flat chain. Here, $\Theta(x)$ is the step function. This simplification allows us to write a linearized (26), without accounting for chemical currents along the $x$ direction. We note that fluctuations along the body axis ($y$ direction) mean that $S_{i}$ is random quantity at each time step. The deterministic solution to this is
@@
We next probe the early time regime of $t<<\tau_{f}$; this corresponds to the approximation of a flat surface (c.f. Fig. 1(d)). We then plug (27) into (25). We then use the small argument expansion for $\cosh(z)\approx 1+\frac{z^{2}}{2}$, with $z=\frac{\chi_{r}S_{i}t}{2}$. This gives us
$\dot{x}_{i}=F_{i}(t)\approx v_{s}\Big(1-\frac{b^{6}}{2\tau_{f}^{2}}S_{i}^{2}t^{2}\Big)+F^{b}_{x,i}+\xi_{t,i}$ (28)
@@
$\langle X\rangle=v_{s}t-\frac{b^{6}}{6\tau_{f}^{2}}\langle S^{2}\rangle t^{3}$ (29)
where $\langle\rangle$ is performed over both the monomers and realizations. For a flat chain (no displacements along propulsion direction), $S_{i}$ is given by the deterministic formula above. In general, Here, $X$ is a site-independent average of the chain position. We obtain the scaling for $(\Delta h)^{2}$ as follows
$W_{h}^{2}$ $\approx\frac{b^{6}}{6\tau_{f}^{2}}\langle\Big(S_{i}^{2}-\langle S\rangle^{2}\Big)^{2}\rangle t^{6}+2D_{t}t$ (30)
where we have used that $\langle F^{b}(t^{\prime})\rangle=0$ and $\langle F^{b}(t^{\prime})F^{b}(t^{\prime})\rangle=\langle F^{b}(t)\rangle\langle F^{b}(t^{\prime})\rangle$. Strictly speaking, the latter will cease to be true when curvature sets in (spring forces then become correlated), but it will suffice for the flat approximation used here. We thus obtain the result that the dynamics is super-ballistic with exponent $\sim 3$ (with an early time diffusive $\sim 0.5$). Here, fluctuations along the body axis minimally accounted for such that $\text{Var}(S)\neq 0$. We thus conclude that minimal roto-translational coupling on a flat interface can predict diffusive to super-ballistic growth transitions but clearly overestimates the super-ballistic exponent to that of a surface with finite curvature.
.4 Early-time approximation to ballistic orientational growth
The deterministic solution on a quasi-flat chain also enables us to rationalize the $\beta_{\theta}\sim 1$ at early times. From (27), let us probe $t<<\tau_{f}$, where we find
$\theta_{i}\approx\frac{1}{\sqrt{2}}(\chi_{r}S_{i}t)[1+\frac{\chi_{r}S_{i}t}{2}]$ (31)
Next, we exploit the symmetries of the orientations along the body axis (c.f. Fig. 3(c)), where we find that $\langle\theta\rangle_{i}=0$ and by extension $\langle(\Delta\theta)^{2}\rangle=\langle(\theta_{i}-\langle\theta\rangle_{i})^{2}\rangle=\langle\theta_{i}^{2}\rangle$. Thus
$\langle(\Delta\theta)^{2}\rangle$ $=\langle\theta_{i}^{2}\rangle=\frac{b^{6}}{2\tau_{f}^{2}}\langle S_{i}^{2}\rangle t^{2}+O(t^{3})+O(t^{4})$ (32)
where the terms $O(t^{3,4})$ are sub-leading for $t<<\tau_{f}$. Thus, a combination of the symmetry of the orientations along the chain ($\theta_{i}=-\theta_{N-i}$) and the transverse fluctuations approximation of the previous section gives us $\beta_{\theta}\sim 1$ at early times.
.5 Curvature of chain
@@
where $h_{i}$ is the height function of the chain, evaluated at each monomer location, measured from the vertical line connecting the edge monomers. The average is performed across the monomers in the chain, giving a scalar value. In our case, the inner bracket is simply reduced to $\partial_{y}\left[{\partial_{y}x_{i}(y)}/{\sqrt{1+|\partial_{y}x_{i}|^{2}}}\right]$. For Fig. 1, we take the averaged $|\kappa|$ in the steady state. Further, we note that $\kappa$ $\rightarrow 0$ as $N\rightarrow\infty$ [29]. For a circular interface, $\kappa$ would be scale as $1/R$ ($R$ the radius of the circle), whereas for a flat interface $\kappa$ would be constant. Thus, this variation the $\kappa$ profile can be used as proxy definition for the C-shape.
Figure 8: Scaling of orientational fluctuations in locally flat region. A $W_{\theta}$ for different locally flat lengths $N^{\prime}$, with the dynamic $\beta_{\theta}^{\prime}$ labelled (fitting window identical to Fig. 3). The same scaling is plotted in the inset. B Roughness scaling for $W_{\theta}^{\text{SS}}(N^{\prime})$. Here, $N=312$ and $\tau_{f}=0.2$, $\Lambda=0.2$, $\tilde{\Lambda}=0.01$.
@@
.6 Scaling of $W_{\theta}$ in the locally flat regime